# How simple can you go? An off-the-shelf transformer approach to molecular dynamics

### v2

@@ 4.3 Q1: Benchmark Results
In Table 1, we present the mean absolute error (MAE) of force predictions for MD-ET, SpookyNet [13], and PaiNN [21].
@@
MD-ET achieves the lowest MAE among the compared models, outperforming both SpookyNet and PaiNN. We observe that SpookyNet’s results do not separate conformers between training and testing, which may overestimate generalization performance. This improvement over equivariant models like PaiNN and SpookyNet suggests that approximate equivariance, as utilized in MD-ET, does not hinder performance on such benchmarks. However, a low test error does not necessarily imply a low MD simulation error or a stable trajectory, as we will demonstrate below.
MD17-10k: molecule specific results
@@
Table 2: Performance comparison on the MD17 benchmark on stability (ps) with a maximum at 300 ps, force MAE (meV Å-1) and FPS. Higher is better is denoted ()($\uparrow$)( ↑ ) and vice versa. Best average results in bold. Results marked with a indicate that the molecule was included in the pretrain train split.
@@
Metric Molecule DeepPot-SE [54] SchNet [17] DimeNet [55] PaiNN [21] SphereNet [56] ForceNet [33] GemNet-dT [34] NequIP [57] MD-ET (pretrained) (ours) MD-ET (finetuned) (ours)
Stability ($\uparrow$) Aspirin 9 26 54 159 141 182 192 300 154 300
Ethanol 300 247 26 86 33 300 300 300 300 300
Naphthalene 246 18 85 300 6 300 25 300 256 300
Salicylic Acid 300 300 73 281 36 1 94 300 213 300
Average 213.8 147.8 60.0 206.5 54 195.8 152.8 300 230.8 300
MAE ($\downarrow$) Aspirin 21.0 35.6 10.0 9.2 3.4 22.1 5.1 2.3 4.2 4.2
Ethanol 8.9 16.8 4.2 5.0 1.7 14.9 1.7 1.3 2.5 $1.0^{*}$
Naphthalene 13.4 22.5 5.7 3.8 1.5 9.9 1.9 1.1 4.0 2.3
Salicylic Acid 14.9 26.3 9.6 6.5 2.6 12.8 4.0 1.6 3.3 2.8
Average 14.6 25.3 7.4 6.1 2.3 15.0 3.2 1.6 3.5 2.6
FPS ($\uparrow$) Aspirin 88.0 108.9 20.6 85.8 17.5 137.3 56.8 8.4 159.8 159.8
Ethanol 101.0 112.6 21.4 87.3 30.5 141.1 54.3 8.9 148.1 148.1
Naphthalene 109.3 110.9 19.1 92.8 18.3 140.2 53.5 8.2 150.2 150.2
Salicylic Acid 94.6 111.7 19.4 90.5 21.4 143.2 52.4 8.4 147.3 147.3
Average 98.2 111.0 20.1 89.1 21.9 140.5 54.3 8.5 151.4 151.4
The MD17 dataset [16] consists of molecular dynamics trajectories for small organic molecules and has been proposed as a benchmark by Fu et al. [58], challenging models to not only predict a low force error but also measures stability, i.e., how long a simulation experiences no catastrophic failure. The third criterion is speed, i.e., how many times per second the model can be evaluated. Table 2 compares MD-ET with several state-of-the-art models, including NequIP [57], GemNet-dT [34]. In addition to training ET-MD directly on MD17, we also report fine-tuning results after 2000 fine-tuning steps (see Appendix A.6 for details). Since there is some overlap with the pretraining set, fine-tuning results for ethanol (marked with a *), have only limited significance.
@@
MD-ET achieves force MAEs similar to those of NequIP and GemNet-dT (which is also not energy-conserving). The only other model that violates energy conservation is ForceNet at a much higher MAE. While NequIP attains slightly lower MAEs, MD-ET excels in computational efficiency, processing over 150 frames per second (FPS) beating all other models. MD-ET also reaches the maximum simulation stability score across all molecules. While a 300 ps NVT simulation is not sufficient to assess the long-term stability of MD trajectories, it can still be used to compare the stability of different models.
@@
For molecular dynamics (MD) simulations where precise electron correlation is necessary to accurately model potential energy surfaces and critical interactions, such as transition states, Density Functional Theory (DFT) reference data may not always be sufficiently accurate. The CCSD(T) level of theory is often regarded as the “gold standard” of quantum chemistry and is thus a suitable choice for these cases. Nonetheless, the substantial computational cost of CCSD(T) renders it impractical for large-scale or long-duration simulations. Transfer learning from lower-level theories, such as DFT, presents an effective strategy to approximate CCSD(T) using the expensive CCSD(T) samples efficiently. We use the MD17@CCSD(T) dataset [20], which recalculates trajectories for molecules in MD17 using CCSD(T) to test this approach using MD-ET. Although MD-ET was pretrained on the PBE0 level, Table 3 demonstrates that MD-ET can effectively capture higher-level interactions through fine-tuning with a limited number of CCSD(T) samples. When fine-tuned on 1,000 CCSD(T)-level samples, MD-ET achieves lower mean absolute error (MAE) than all competing models for three of the five structures. For toluene, its MAE is comparable to GemNet-T, while for benzene, it exhibits the highest error among the benchmarked models. Furthermore, when we reduce the fine-tuning dataset by an order of magnitude MD-ET still reaches accuracy comparable to models trained on the full dataset.
Ko2020: molecule specific results
Ko et al. [53] introduce a molecular dataset consisting of molecular (e.g. $\text{C}_{10}\text{H}_{2}$ and $\text{C}_{10}\text{H}^{+}$) and metallic systems (e.g., $\text{Ag}_{3}^{+/-}$), which are typically challenging for conventional MLFFs. Table 4 compares the force root mean squared errors (RMSE) of MD-ET with 4G-BPNN [53] and SpookyNet [13] on selected systems from the Ko2020 dataset.
Table 4: RMSEs of atomic forces (meV Å-1) for different molecular systems, comparing 4G-BPNN, SpookyNet, and our work. The values for 4G-BPNN are taken from [53]. Best results in bold.
@@
For the C10H2/C10H+${}_{3}^{+}$ and Ag+/${}_{3}^{+/-}$ systems, MD-ET outperforms both 4G-BPNN and SpookyNet. This demonstrates MD-ET’s capability to model smaller ionic and metallic systems. However, on the Na8/Cl+${}_{8}^{+}$ MD-ET’s performance is slightly worse and fails to converge on Au2–MgO systems.
4.4 Q2: Equivariance Evaluation
Figure 1: Equivariance Evaluation. Top Left: Average equivariance error $E_{\text{eq}}$ for 2048 randomly sampled QCML train and test structures using different frame-averaging sample sizes $n$. Top Right: Equivariance error (without frame averaging) over the $\mathrm{SO(3)}$ group across a range of alkanes and cumulenes. Bottom Left: Approximate numerical noise level over the $\mathrm{SO(3)}$ group for alkanes and cumulenes. Bottom Right: Relative equivariance error for alkanes and cumulenes.
Our model, MD-ET, is only approximately equivariant. Although MD-ET’s performs well on benchmarks, we examine the potential influence of approximate equivariance on MD simulations more closely. To quantify how (non-)equivariant our model is, we define the equivariance error as the expected Euclidean distance between the model’s prediction on a randomly rotated input and the expected prediction averaged over all rotations in the symmetry group,
$E_{\text{eq}}({D},\hat{f}_{\theta})=\underset{\begin{subarray}{c}x\sim{D}\\ \mathcal{S}\sim\text{SO}(3)\end{subarray}}{\mathbb{E}}\left[\left\|\underset{\mathcal{R}\sim\text{SO}(3)}{\mathbb{E}}\left[\mathcal{R}^{\top}\hat{f}_{\theta}(\mathcal{R}{x})\right]-\mathcal{S}^{\top}\hat{f}_{\theta}(\mathcal{S}{x})\right\|_{2}\right],$ (5)
where $x\sim{D}$ are the 3D positions sampled from their distribution $D$, $\mathcal{R},\mathcal{S}\sim\text{SO}(3)$ are rotation matrices, and $\hat{f}_{\theta}$ is a trained predictor. For a perfectly equivariant model, $E_{\text{eq}}=0$ (proof in Appendix A.4).
In practice, models use floating point arithmetic with finite precision and numerical inaccuracies accumulate during evaluation. For this reason, equivariant models do not reach $E_{\text{eq}}=0$. To better assess MD-ET’s deviation from an equivariant architecture, we evaluate $E_{\text{eq}}$ using different amounts of frame-averaging (see section 4.2). Concretely, this means averaging the model predictions over $n$ randomly sampled rotations $\mathcal{R}_{i}\in\mathrm{SO(3)}$,
$\hat{f}_{\theta}^{\prime}(x)=\int_{\mathrm{SO(3)}}\mathcal{R}^{\top}\hat{f}_{\theta}(\mathcal{R}x)\,\mathrm{d}\mu(\mathcal{R})\approx\frac{1}{n}\sum_{i=1}^{n}\mathcal{R}_{i}^{\top}\hat{f}_{\theta}(\mathcal{R}_{i}x).$ (6)
The limit, $\lim_{n\to\infty}E_{\text{eq}}(\mathcal{D},f^{\prime})$, converges to zero for any $f_{\theta}^{\prime}$ (proof in Appendix A.4). When empirically evaluating $E_{\text{eq}}(\mathcal{D},\hat{f}_{\theta}^{\prime})$ for high values of $n$, we can thus attribute remaining errors to numerical inaccuracies during evaluation. We can thus grasp the numerical limit of $\mathrm{SO}(3)$ equivariance and how close our model has approached it. In practice we approximate $\lim_{n\to\infty}E_{\text{eq}}(\mathcal{D},f^{\prime})$ by evaluating $E_{\text{eq}}$ for multiple frame-averaging sizes $n$ and fitting an exponential function to extrapolate to the asymptotic value for $n\to\infty$.
We evaluate MD-ET’s average $E_{\text{eq}}$ across 2048 randomly drawn QCML train and test samples (see Figure 1, top left). The results show that without frame-averaging MD-ET reaches an $E_{\text{eq}}$ of around $0.06$ kcal/mol/Å (top left), which is one order of magnitude below our test set loss and two to three orders of magnitude below typical force magnitudes. Frame-averaging exponentially decreases the equivariance error towards a value below $0.02$ kcal/mol/Å (bottom right). $E_{\text{eq}}$ is higher on the test set, which suggests that learned equivariance might not be retained further from the training distribution. To investigate this possibility, we calculate $E_{\text{eq}}$ for a number of alkanes (purple) and cumulenes (blue) of increasing size: Structures with more than 8 carbons are out-of-distribution (see Figure 1, top right; for visualization, see Appendix B). For each structure, we approximate the numerical noise level (see Figure 1, bottom left) using MD-ET and an equivariant SpookyNet. We also examine the ratio between the noise level and $E_{\text{eq}}$ without frame-averaging (see Figure 1, bottom right) to evaluate how closely MD-ET approaches the practical limit for SO(3) equivariance across structures.
Results demonstrate that MD-ET’s approximation of SO(3) equivariance measured relative to the noise level varies little between structures, even in the extrapolative regime. On the other hand, numerical noise varies significantly between structures and tends to increase with structure size. This observation extends to the equivariant SpookyNet architecture.
For in-distribution data, we find that MD-ET learns to approximate equivariance well, approaching the limit imposed by the numerical accuracy to an order of magnitude. Given that we use only one additional data augmentation per sample during training, learning approximate equivariance appears cheap and effective. The achieved equivariance error is unlikely to impact force prediction accuracy since it is two orders of magnitude lower than typical molecular forces for most structures. Four frame-averaging samples can further reduce the equivariance error by more than 50 %. In the extrapolative regime, we observe an increase in numerical noise and larger differences in noise levels between structures. While problems with out-of-distribution generalization are widely studied, deteriorating numerical stability in the extrapolative regime is a distinct phenomenon and has a significant effect on an MLFF’s prediction quality. In contrast, learned equivariance appears to generalize well when noise levels are taken into account. We thus conclude that learning approximate equivariance is unlikely to impede a model’s accuracy, while numerical instability–which is a concern for all ML models–could.
4.5 Q3: Evaluation of Energy Conservation
Figure 2: Total energy variation over time for NVE simulations of linear alkanes CnH2n+2 ($n=1\dots 8$). For reference, we compare to MD simulations with an energy-conserving model (SpookyNet [13]) trained on the same data as MD-ET. A linear fit is plotted above the raw data to help visualize energy drift.
Since MD-ET directly predicts forces without enforcing energy conservation, we aim to assess whether the predicted forces are approximately conservative in addition to being approximately equivariant. Although MD-ET achieves stable MD simulations on the MD17 benchmark (see Table 2), this alone is not sufficient to demonstrate approximate energy conservation. This is because the dynamics sample the canonical (NVT) ensemble, i.e., the aim is to keep the temperature of the simulated system constant. To achieve this, a thermostat constantly introduces or removes energy, which prevents testing energy conservation. We therefore also perform MD simulations sampling the microcanonical (NVE) ensemble, where total energy is a conserved quantity. This enables quantification of how strongly a simulation violates energy conservation by monitoring the variation of total energy over time. Note that even when using conservative FFs, the total energy typically fluctuates slightly due to numerical noise and discretization errors when integrating the equations of motion. This is acceptable as long as these fluctuations are small and stay centered around zero, i.e., do not introduce energy drift. We find that NVE simulations with MD-ET are approximately energy-conserving for small systems (fewer than around 15 atoms), but for larger systems, the total energy tends to increase linearly with time (see Figure 2). The additional energy causes increasingly large structural fluctuations until the simulation invariably becomes unstable after a sufficiently large number of time steps. For many MD experiments MD-ET’s utility is thus questionable. Nevertheless, our results are surprising: To our knowledge no other approximately energy-conserving MD model has been shown to be approximately energy-conserving and stable in short NVE simulations [29].
4.6 Q4: MD Simulation results
Figure 3: Left column: Dihedral angle of cumulenes of length 6 to 9 over an NVT simulation of 30 ps. Right column: Instantaneous temperature over time. The inset in the first two rows shows the flat structure of the first cumulene and the perpendicular structure in the second one. The dihedral angle $\omega$ is denoted in red, encompassing the four hydrogen atoms.

### v3

@@ 4.1 Unconstrained models are competitive on common MD benchmarks
In Table 1, we present the mean absolute error (MAE) of force predictions for MD-ET, SpookyNet [72], and PaiNN [67] on the QCML evaluation set at the end of pretraining.
@@
MD-ET achieves the lowest MAE among the compared models, outperforming both SpookyNet and PaiNN. This improvement over equivariant models like PaiNN and SpookyNet suggests that approximate equivariance, as utilized in MD-ET, does not hinder performance on such benchmarks. However, a low test error does not necessarily imply a low MD simulation error or a stable trajectory, as we will demonstrate below.
MD17-10k
@@
Table 2: Performance on the MD17-10k benchmark, comparing stability (ps) with a maximum at 300 ps, force MAE (meV Å-1), the distribution of interatomic distances, $h(r)$, and FPS. Higher is better is denoted $(\uparrow)$ and vice versa. Best average results in bold. Results marked with a indicate that the molecule was included in the pretrain train split. FPS benchmarks conducted on a V100 GPU, except where marked ∗∗, there conducted on an A100 GPU.
@@
Metric Molecule DeepPot-SE [81] SchNet [66] DimeNet [28] PaiNN [67] SphereNet [45] ForceNet [34] GemNet-dT [27] NequIP [5] MD-ET (trained directly) (ours) Orb [54] (finetuned) MD-ET (finetuned) (ours)
Stability ($\uparrow$) Aspirin 9 26 54 159 141 182 192 300 124 300 300
Ethanol 300 247 26 86 33 300 300 300 300 300 300
Naphthalene 246 18 85 300 6 300 25 300 300 300 300
Salicylic Acid 300 300 73 281 36 1 94 300 256 300 300
Average 213.8 147.8 60.0 206.5 54 195.8 152.8 300 245 300 300
MAE ($\downarrow$) Aspirin 21.0 35.6 10.0 9.2 3.4 22.1 5.1 2.3 4.2 2.4 4.2
Ethanol 8.9 16.8 4.2 5.0 1.7 14.9 1.7 1.7 1.3 2.5 $1.0^{*}$
Naphthalene 13.4 22.5 5.7 3.8 1.5 9.9 1.9 1.1 4.0 1.0 2.3
Salicylic Acid 14.9 26.3 9.6 6.5 2.6 12.8 4.0 1.6 3.3 1.8 2.8
Average 14.6 25.3 7.4 6.1 2.3 15.0 3.2 1.6 3.5 1.7 2.6
h(r) ($\downarrow$) Aspirin 0.65 0.36 0.04 0.04 0.03 0.56 0.04 0.02 0.03 0.03 0.02
Ethanol 0.09 0.21 0.15 0.15 0.13 0.86 0.09 0.08 0.03 0.09 0.05
Naphthalene 0.11 0.09 0.10 0.13 0.14 1.02 0.12 0.12 0.15 0.12 0.07
Salicylic Acid 0.03 0.03 0.06 0.03 0.06 0.35 0.07 0.03 0.3 0.03 0.03
Average 0.22 0.17 0.088 0.088 0.090 0.70 0.080 0.063 0.125 0.068 0.04
FPS ($\uparrow$) Aspirin 88.0 108.9 20.6 85.8 17.5 137.3 56.8 8.4 159.8 68.6** 159.8
Ethanol 101.0 112.6 21.4 87.3 30.5 141.1 54.3 8.9 148.1 69.2** 148.1
Naphthalene 109.3 110.9 19.1 92.8 18.3 140.2 53.5 8.2 150.2 68.5** 150.2
Salicylic Acid 94.6 111.7 19.4 90.5 21.4 143.2 52.4 8.4 147.3 69.4** 147.3
Average 98.2 111.0 20.1 89.1 21.9 140.5 54.3 8.5 151.4 68.9** 151.4
The MD17 dataset [15] consists of MD trajectories for small organic molecules and has been proposed as a benchmark by Fu et al. [25]. Models are evaluated not only for their prediction accuracy (MAE) but also for their simulation stability and faithfulness. Stability is measured by the time until a simulation becomes unstable (capped at 300 ps), and faithfulness is measured by $h(r)$, the distribution of interatomic distances [25]. Inference speed, which is critical for MD simulations, is also compared (see Appendix A.6 for evaluation details) in frames per second (FPS).
Table 2 compares MD-ET with several state-of-the-art models. In addition to fine-tuning ET-MD for 2000 steps, which we see as the default way to use a large-scale pretrained model, we also report results for training the model directly on MD17. Since there is some overlap with the pretraining set, fine-tuning results for ethanol (marked with a *), have only limited significance. We also include fine-tuning results for Orb-v2 [54], which was released recently and is a popular unconstrained model.
@@
MD-ET compares favorably to competing models in 3 out of 4 metrics. MD-ET’s MAE score is slightly worse than that of several other models but it produces the most faithful NVT MDs according to the h(r) score (Fu et al. [25]). MD-ET also reaches the maximum simulation stability score across all molecules. Although a 300 ps NVT simulation is not sufficient to assess the long-term stability of MD trajectories, it can still be used to compare the stability of different models.
@@
For MD simulations where precise electron correlation is necessary to accurately model potential energy surfaces and critical interactions, density functional theory (DFT) reference data may not always be sufficiently accurate. Coupled cluster (CC) including single and double excitations with the perturbative triples correction (CCSD(T)) is often regarded as the “gold standard” of quantum chemistry and is thus a suitable choice for these cases. However, the substantial computational cost of CCSD(T) renders it impractical for large-scale or long-duration simulations. Transfer learning from lower-level theories, such as DFT, presents an effective strategy to approximate CCSD(T) using the expensive CCSD(T) samples efficiently [37]. We use the MD17@CCSD(T) dataset [14], which recalculates trajectories for molecules in MD17 using CCSD(T) to test this approach using MD-ET. Although MD-ET was pretrained on the PBE0 level, Table 3 demonstrates that MD-ET can effectively capture higher-level interactions through fine-tuning with a limited number of CCSD(T) samples. When fine-tuned on 1,000 CCSD(T)-level samples, MD-ET achieves lower mean absolute error (MAE) than all competing models for three of the five structures. For toluene, its MAE is comparable to GemNet-T, while for benzene, it exhibits the highest error among the benchmarked models. Furthermore, when we reduce the fine-tuning dataset by an order of magnitude MD-ET still reaches accuracy comparable to models trained on the full dataset (see Appendix A.6.2 for experimental details).
Ko2020
Table 4: RMSEs of atomic forces (meV Å-1) for different molecular systems, comparing 4G-BPNN, SpookyNet, and MD-ET. The values for 4G-BPNN are taken from [39]. Best results in bold.
@@
Ko et al. [39] introduce a molecular dataset consisting of molecular (e.g. $\text{C}_{10}\text{H}_{2}$ and $\text{C}_{10}\text{H}^{+}$) and metallic systems (e.g., $\text{Ag}_{3}^{+/-}$), which are typically challenging for conventional MLFFs. Table 4 compares the force root mean squared errors (RMSE) of MD-ET with a 4th generation Behler–Parinello neural network, 4G-BPNN (high-dimensional neural network potential, 4G-HDNNP) [39] and SpookyNet [72] on selected systems from the Ko2020 dataset.